OpenStudy (tpaulus):

Where does the formula vy = .5gt come from?

OpenStudy (tpaulus):

\[{v_y} = \frac{1}{2}gt\]

OpenStudy (anonymous):

It's generally derived with calculus - want to have a go at it? ^_^

OpenStudy (tpaulus):

I'm in AP Calc AB, so I'm down...

OpenStudy (anonymous):

It'll be cake then ^_^ no lie. So, you start with your definition of acceleration, \[a=\frac{\mathrm{d}v}{\mathrm{d}t}\] that multiply by dt and you get \[a \mathrm{d} t =\mathrm{d}v\] So we make that the indefinite integral \[\int a \mathrm{d} t =\int\mathrm{d}v\] What does that give you?

OpenStudy (anonymous):

a is a constant in this scenario, btw, so yo can say \[a \int \mathrm{d} t =\int\mathrm{d}v\]

OpenStudy (anonymous):

Let's also actually just put boundaries on it, to make the constant behave. Time starts at zero, and your velocity starts at some value v0 \[a\int_0^t \mathrm{d} t =\int_{v_0}^v\mathrm{d}v\]

OpenStudy (anonymous):

How would you evaluate \[\int_0^x 1 \mathrm{d}x\]?

OpenStudy (tpaulus):

∫dx = x right?

OpenStudy (anonymous):

Exactly! It's just x (plus a constant) evaluated from x to zero, right? \[\int_0^x 1 \mathrm{d}x = (x+C)\big|_0^x\] \[ = (x+C)-(0+C)=x\] \[\int_0^x 1 \mathrm{d}x = x\] Similarly, looking at the left hand side of the original equation \[a\int_0^t \mathrm{d}t = a \Big(t+C\Big) \Big|_0^t=a(t)=at\] \[a\int_0^t \mathrm{d}t=at\] Now looking at the right hand side \[ \int_{v_0}^v \mathrm{d}v = (v+C)\Big|_{v_0}^v=(v+C)-(v_0-C)=v-v_0\] \[\int_{v_0}^v \mathrm{d}v=v-v_0\] together they give the equation for the velocity of a particle given its acceleration and time \[a\int_0^t \mathrm{d}t=\int_{v_0}^v \mathrm{d}v\] \[at=v-v_0\] \[v=at+v_0\] magic!

OpenStudy (tpaulus):

Thanks, my Physics Teacher is going to be amazed!

OpenStudy (anonymous):

^_^ Can you see where the next step comes is? this is just for the velocity, not the position. Do you know the definition formula for velocity (like the one for acceleration above?)

OpenStudy (tpaulus):

v= d/t

OpenStudy (anonymous):

Correct - do you know the differential form? (and if not, the form with Deltas?)

OpenStudy (anonymous):

like, \[a=\frac{\mathrm{d}v}{\mathrm{d}t}\] whereas \[v=?\]

OpenStudy (tpaulus):

\[a = \frac{ dd }{ dt }\]

OpenStudy (anonymous):

Close! ^_^ To prevent having to write dd (which might be confusing) we'll use dx - just some change in x (or it could be in the y axis or the z axis or whatever - it's just a change in **position**) So we have \[v=\frac{\mathrm{d}x}{\mathrm{d}t}\] Like before, we can separate that into the expression \[v\mathrm{d}t=\mathrm{d}x\] And then into the integral \[\int_0^t v \mathrm{d}t=\int_{x_0}^x \mathrm{d}x\]

OpenStudy (tpaulus):

\[v=\frac{ dd }{ dt }\]

OpenStudy (anonymous):

So unlike before with a, v is NOT a constant. It's the expression that we found above from our first integral! \[v=at+v_0\] So we plug that into our left hand side \[\int_0^t v \mathrm{d}t=\int_0^t (at+v_0)\mathrm{d}t\]

OpenStudy (anonymous):

can you see what that expression gives?

OpenStudy (anonymous):

both a and v0 are constants, and the variable you're integrating over is t

OpenStudy (tpaulus):

\[v=\frac{ 1 }{ 2 }at^2 + v _{0}t\]

OpenStudy (anonymous):

your evaluating is perfect! It's not v that's equal to that, however, it's the integral \[\int_0^t(at+v_0)\mathrm{d}t=\frac{1}{2}at^2+v_0t\] Which is the left hand side of the original equation. \[\int_0^t(at+v_0)\mathrm{d}t = \int_{x_0}^x\mathrm{d}x\] So \[\frac{1}{2}at^2+v_0t = \int_{x_0}^x\mathrm{d}x\] What's the right hand side then?

OpenStudy (anonymous):

Also, I just noticed that your equation is \[v_y=.5gt\] I had misread that and thought it was position, not velocity. Is there a specific problem that gives that velocity as an answer? (this derivation is still useful to know (and fun to do!), since it's the basis for all kinematics, but I'm unsure where you got your equation from now).

OpenStudy (tpaulus):

It's for a Physics Lab, where we have to calculate the initial velocity of the y-component of a thrown ball. My physics teacher gave us the formula, but we need to prove it.

OpenStudy (tpaulus):

We measured the time using a timer.

OpenStudy (tpaulus):

If you take the derivative of the position function, you would get the velocity function.

OpenStudy (anonymous):

That statement is correct. Did you throw the ball straight with the horizontal?

OpenStudy (tpaulus):

No, it was thrown at an angle, by calculating the x and y initial velocities, i used the formula \[\theta = \left| \tan^{-1} \frac{ v _{y} }{ v _{x}} \right|\] and calculated an angle of about 65 degeres

OpenStudy (anonymous):

65º above the horizontal? Did yo measure either the final velocity of the ball, or the distance it was thrown?

OpenStudy (tpaulus):

final distance

OpenStudy (tpaulus):

Here's my draft... If you want to read/skim it... (It still has a lot of typos)

OpenStudy (anonymous):

You seem to already have the initial y velocity... (I'm sorry, just trying to figure out what to do).

OpenStudy (tpaulus):

Yes, I have the formula, but I need to provide a proof for the formula

OpenStudy (tpaulus):

The proof I tired was wrong, see page 3

OpenStudy (anonymous):

That statement as a value for initial velocity doesn't make sense to me - the initial velocity wouldn't be a function of anything. It's just a value.

OpenStudy (tpaulus):

Your first equation, that resulted in v=at+v0; is what I'm looking for. The ball reaches it's max at 1/2 t, and it's initial velocity is 0m/s, because it's thrown.

OpenStudy (anonymous):

your second statement is not true. If its initial velocity in the y direction was zero, it would only fall and never travel upwards.

OpenStudy (tpaulus):

It's being accelerated though

OpenStudy (anonymous):

Only downwards by the force of gravity.

OpenStudy (tpaulus):

Hmmm...

OpenStudy (anonymous):

And you already used an initial velocity to calculate the angle it was being thrown. I'm sorry, I just don't know what you're looking for, or what your teacher gave. Did your teacher really say that the initial velocity in the y direction was \[v_{yi}=\frac{1}{2}gt\] ? Or that the velocity in the y direction is equal to that - those are very different statements.

OpenStudy (tpaulus):

Initial Y velocity, sorry about that

OpenStudy (anonymous):

That's what you had said before, and makes very little sense to me at the moment.. hmm... Initial velocity exists at t=0, so by definition of it being an initial value that expression is zero.

OpenStudy (anonymous):

ie, \[v=v_i+at\] by definition, an initial velocity is when t=0 \[v=v_i +a(0)\] \[v=v_i \quad @ t=0\]

OpenStudy (tpaulus):

The Instructions say, "Calculate initial vertical velocity. This is the acceleration due to gravity, 9.81 meters per second squared or 32 feet per second squared, times half the total time the ball was in the air."

OpenStudy (anonymous):

OH! It's not a function of t! t is a delta t. Dealio. k, 1 sec.

OpenStudy (tpaulus):

Glad that that helped

OpenStudy (anonymous):

Definitely was in the wrong headspace. Okay, so yeah. You basically already did it up top ^_^ Your derivation in the lab didn't work because you used \[v=d/t\] which gives a constant velocity over some distance. The ball is being acceleration. So you can either use the integration method from up top, or take the derivative of the position function wrt to t \[x(t)=v_0t+\frac{1}{2}at^2\] \[\frac{\mathrm{d}x}{\mathrm{d}t}=v_0 + at\] \[v(t)=v_0+at\] Now the fun part!! So, looking at the graph  What's the y component of the velocity at the top of the parabola?

OpenStudy (tpaulus):

vy = 0

OpenStudy (anonymous):

exactly, so using \[ t_{Tot} = \text{total time}\] \[t_{top} = \frac{1}{2} t_{Tot}\] We have our function \[v_y(t)=v_{0y}-gt\] evaluated at time t_top, v_y=0 \[v_y(t_{top})=0\] \[0=v_{0y}-g(t_{top})\] \[v_{0y}=g(t_{top})\] \[v_{0y}=\frac{1}{2}g \ t_{Tot}\]

OpenStudy (anonymous):

I'm sorry I made that so complicated before :P Brain death!

OpenStudy (tpaulus):

No Problem, Thanks alot!!

OpenStudy (anonymous):

Very welcome. Good luck with the write up! ^_^

OpenStudy (tpaulus):

Thanks!